OpenStudy (anonymous):

i have 22dimes and nickles the value of the coins is 1.85. how many of each coin i have

OpenStudy (amistre64):

im gonna guess, and this needs to be dbl chked amount dimes = 22(1.85 - .05)/(.10-.05)

OpenStudy (amistre64):

792 aint it lol

OpenStudy (anonymous):

yeah thats what i got

OpenStudy (amistre64):

n + d = 22 n(.05) + d(.10) = 1.85 ahh, i see it

OpenStudy (amistre64):

n + d = 22 n(.05) + d(.10) = 1.85 n(-.05) + d(-.05) = 22(-.05) n(.05) + d(.10) = 1.85 ---------------------- d(.10-.05) = 1.85 - 22(.05) d = (1.85-22(.05))/(.10-.05)

OpenStudy (amistre64):

15 dimes and the rest are nickels

OpenStudy (anonymous):

thank you like a thousand times

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

convert pennies. {d + n = 22, 10 d + 5 n =185} {d = 15, n = 7}

OpenStudy (anonymous):

convert to pennies.